Herb packs fish in 500 gram cans. He is paid $8.25 per hour plus $0.18 per kilogram in excess of 500 kg in 7.5 hour shift. How much will he earn per day if he packs 250 cans per hour?
Found 2 solutions by mananth, MathTherapy:
Answer by mananth(16946)

Herb packs fish in 500 gram cans. He is paid $8.25 per hour plus $0.18 per kilogram in excess of 500 kg in 7.5 hour shift. How much will he earn per day if he packs 250 cans per hour?
In 7.5 hours herb packs
250 cans per hour each 500 grams (0.5 kg)
Total packed in a day = 250 * 0.5 *7.5 = 937.5
Wages on hourly basis 8.25 *7.5 = 61.88
Extra packed including partially packed 937.5-500= 437.50 kg
Wage due to excess packing = 0.18 *437.50 = 78.75
Total earningin a day = 61.88 +78.75 = $140.63

Answer by MathTherapy(10552)

Herb packs fish in 500 gram cans. He is paid $8.25 per hour plus $0.18 per kilogram in excess of 500 kg in 7.5 hour shift. How much will he earn per day if he packs 250 cans per hour?

Number of hours worked in a day: 7.5
Hourly-salary, per 7.5-hour day: 7.5(8.25) = $61.88

Number of cans packed in a 7.5-hour day, at 250 cans per hour: 7.5(250) = 1,875
Weight of 1,875 cans packed, in a 7.5-hour day: 1,875(500) = 937,500 gms, or 937.5 kg
Amount of kilograms, packed per day, in excess of 500 kg: 937.5 - 500 = 437.5
Bonus, @ $0.18 per kg in excess of 500 kg = .18(437.5) = $78.75

Total earnings per 7.5-hour day, including bonus: $61.88 + 78.75 = $140.63.